Brief Exercise 24-8

Answer each of the questions in the following unrelated situations.

(a) The current ratio of a company is 5:1 and its acid-test ratio is 1:1. If the inventories and prepaid items amount to $507,000, what is the amount of current liabilities?

Current Liabilities $


(b) A company had an average inventory last year of $203,000 and its inventory turnover was 5. If sales volume and unit cost remain the same this year as last and inventory turnover is 8 this year, what will average inventory have to be during the current year? (Round answer to 0 decimal places, e.g. 125.)

Average Inventory $


(c) A company has current assets of $86,000 (of which $39,000 is inventory and prepaid items) and current liabilities of $39,000. What is the current ratio? What is the acid-test ratio? If the company borrows $16,000 cash from a bank on a 120-day loan, what will its current ratio be? What will the acid-test ratio be? (Round answers to 2 decimal places, e.g. 2.50.)

Current Ratio

:1
Acid Test Ratio

:1
New Current Ratio

:1
New Acid Test Ratio

:1


(d) A company has current assets of $623,000 and current liabilities of $220,000. The board of directors declares a cash dividend of $189,000. What is the current ratio after the declaration but before payment? What is the current ratio after the payment of the dividend? (Round answers to 2 decimal places, e.g. 2.50.)

Current ratio after the declaration but before payment

:1
Current ratio after the payment of the dividend

:1
Answer #1

(a)
Current Ratio = Current Assets / Current Liabilities
5 x Current Liabilties = Current Assets

Acid Test Ratio = (Current Assets - Inventory and prepaid items) / Current Liabilities
Current Liabilities = 5 x Current Liabilities - $507000
Current Liabilities = $126750

(b)
Inventory Turnover = Cost of Goods Sold / Average Invetntory
5 = Cost of Goods Sold / $203000
Cost of Goods sold = $1015000

Average Inventory = $1015000 / 8 = $126875

(c)
Current Ratio = $86000 / $39000 = 2.21
Acid Test Ratio = ($86000 - 39000) / 39000 = 1.21

Current Ratio = (86000+16000)/(39000+16000) = 1.85
Acid Test Ratio = (86000+16000-39000)/(39000+16000) = 1.15

(d)
Current Ratio = $623000 /(220000+189000) = 1.52

After payment
Current Ratio = (623000-189000) / 220000 = 1.97